Thursday, November 1, 2012

Day 9, November 2, 2012 -- Closer to closing

The cryostat team (Kyle, Jeff, and myself) managed to get the cryostat a fair amount of the way towards being closed up and on the vacuum pump. We removed the instrument after yesterday's aligning, Jeff installed the HWP system and fully tested it, Kyle and I finished the instrument's thermal and electrical connections, we re-installed the optical filters in the optics stack at 4K and 77K, re-installed the instrument for the final time into the cryostat, installed SQUIDs, and connected detectors to their SQUID boards. Whew! Tomorrow, all we have left to do is wiring checks, closing up shells, and installing the window assembly before pumping.

The gondola team was also busy. Michele finished re-assembling the rotator while Britt worked on debugging the temperature channels in the gondola thermometry. In addition, Chappy and Michele reassembled the first star camera (unfortunately named ESC0, because Chappy is a computer scientist at heart who loves zero-indexing) to get it ready for its dry nitrogen purge. Seth was probably also really productive, but it's hard to see progress in software in the high bay on a day to day basis.

Only a few pictures today:

No comments:

Post a Comment

Note: Only a member of this blog may post a comment.